OpenStudy (anonymous):

If you can...Solve the system of equations. 2x - 5y = 3 x - 3y = 1

hartnn (hartnn):

multiply the 2nd equation by 2 what do u get ?

OpenStudy (anonymous):

?

hartnn (hartnn):

the co-efficient of 'x' in 1st equation is 2 but in 2nd equation it is 1 by multiplying the whole 2nd equation by 2, we are trying to make the same co-efficient for 'x' so that we can subtract these equation, and 'x' would get eliminated!

OpenStudy (anonymous):

alright here are the answer choices. (4, 1) (5, 2) (7, 2) (9, 3)

hartnn (hartnn):

sorry, i didn't ask for it. here's your 2nd equation, \(x - 3y = 1\) now, multiply every term by 2 \(2(x - 3y) = 2\times1\) can you simplify that ? (thats what i am asking)

OpenStudy (anonymous):

I don't even know how to do this. That's why i'm here.

hartnn (hartnn):

I am trying to make you learn how to do it, thats why i am here :) i think you can try simplifying left side, \(2(x-3y)\) heard of distribution property ??? \(a(b+c)=ab+ac\)

OpenStudy (anonymous):

ah can I learn tomorrow its almost 2:00 am I just need the answers at this point. ill be back tomorrow can we review then?

hartnn (hartnn):

ofcourse! :) just tag me when you come back ...

OpenStudy (anonymous):

alright could I get the answer do this question tho? c:

OpenStudy (anonymous):

problem* I cant even think straight. :P

hartnn (hartnn):

why are you in a hurry?? you'll get it by yourself when we take it up tomorrow :)

OpenStudy (anonymous):

I already entered the quiz >,<

hartnn (hartnn):

sorry, we are not allowed to help on a graded quiz here on OpenStudy, its just against our Code of Conduct.

OpenStudy (anonymous):

its like a practice quiz its not graded at all.

OpenStudy (anonymous):

We could just do it tomorrow then c:thanks see ya n ill message tomorrow

hartnn (hartnn):

sure, have a good night and sweet dreamzzzz :)